OpenStudy (lucaz):

domain of sqrt(x^2-3) / sqrt(4-x)

OpenStudy (lucaz):

I found sqrt(3)<=x<4

OpenStudy (math&ing001):

How did you find that ?

OpenStudy (lucaz):

x^2-3>=0 and 4-x>0

OpenStudy (math&ing001):

Yeah, so from the first inequality, what would you get ?

OpenStudy (lucaz):

x greater or equal to sqrt of 3

OpenStudy (math&ing001):

yeah your answer is correct.

OpenStudy (math&ing001):

I was confused about it at first because x²>=3 means x>=sqrt(3) or x<=-sqrt(3)

OpenStudy (lucaz):

i think its ok for x less than neg. Sqrt of 3, if you pulg it in..

OpenStudy (lucaz):

both numerator and denominator are defined

OpenStudy (math&ing001):

Yeah that's was I thought, so the domaine would be \[]-\infty,-\sqrt{3} ] \cup [\sqrt{3},4[\]

OpenStudy (lucaz):

ok, thanks for your help

OpenStudy (math&ing001):

Welcome ! p.s: you can see it here : https://www.desmos.com/calculator/z0lnkkgwnj